Myra is paid based on the number of pages she types. She is paid P 5.00 per page plus free meals. For the week, she typed 375 pages. The week being the last week of the month, deductions are made from employee's gross earnings equivalent to 4% of monthly earnings for SSS, P 10.50 for Medicare, she is paying at the rate of P 50.00 per week. Myra's earnings for the current month including the current week is P 8 000.00 Compute for Myra's take-home pay.


1
Expert's answer
2021-05-07T09:07:57-0400

Monthly income of myra including current week "= 8000"


Deduction he have to pay for SSS "= \dfrac{4}{100}\times 8000=320"

Deduction for mrdicare "=10.50"


Per week deduction of myra "= 50"


Myra's take home-pay"=\text{ Monthly income-Total deduction}"

"=8000-(50+320+10.5)=8000-380.5=7619.5"


Hence Myra's take-home pay is "P7619.5"